Jayp:

A rectangle has sides measuring (6x + 4) units and (2x + 11) units. Part A: What is the expression that represents the area of the rectangle? Show your work. (4 points) Part B: What are the degree and classification of the expression obtained in Part A? (3 points) Part C: How does Part A demonstrate the closure property for polynomials? (3 points)

mhanifa:

Part A. Area = w*l, substitute w = 6x + 4, l = 2x + 11 and find the expression Part B. Since each of the multiples are of degree one, the resultant expression will have the degree of 1 + 1. Name the polynomial as you get the result of Part A Lets discuss part C once you complete parts A and B